OpenStudy (anonymous):

What is 5^(2x-2) = 35??

OpenStudy (anonymous):

Well Distribute the 5 and tell me what you get

OpenStudy (anonymous):

x = (3 log(5)+log(7))/(2 log(5)) ?

OpenStudy (anonymous):

Wow I got no clue

OpenStudy (phi):

use this property \[\log(a^b)= b \log(a)\] so your equation becomes (2x-2)log(5) = log(35) or \[ 2x-2= \frac{log(35)}{log(5} \] \[ x= \frac{log(35)}{2log(5)} +1\] you can massage this into all kinds of things using these rules: log(a b)= log(a)+log(b) and the exponent rule. for example 35= 5 x 7 so log(35)= log(5*7)= log(5)+log(7) so you could get \[ x= \frac{log(5)+log(7)}{2log(5)} +\frac{2log(5)}{2log(5)} = \frac{3log(5)+log(7)}{2log(5)} \] or using 2log(5)= log(5^2)= log(25) \[ x= \frac{log(5*25*7)}{log(25)}= \frac{log(875)}{log(25)} \]